OpenStudy (anonymous):

How do you solve the equation? (1)/(3x^2)+(1)/(3x)=(2)/(x)

OpenStudy (anonymous):

\[1/3x ^{2}+1/3x=2/x\]

OpenStudy (anonymous):

Make the denominator the same for the left side and right sides. \[\frac{ 1 }{ 3x^2 }+\frac{ x }{ 3x^2}=\frac{ 6x }{ 3x^2 } \rightarrow \frac{ 1+x }{ 3x^2 }=\frac{ 6x }{ 3x^2 }\] Now multiply both sides by 3x^2 to cancel it out and then solve for x.\[1+x=6x\rightarrow x=\frac{ 1 }{ 5 }\]